aton Electronic Company’s treasurer uses both the capital asset pricing model and the dividend valuation model to compute the cost of common equity (also referred to as the required rate of return for common equity).

Assume: Rf = 8 % Km = 13 % β = 1.7 D1 = $ .90 P0 = $ 20 g = 9 %

a. Compute Ki (required rate of return on common equity based on the capital asset pricing model). (Do not round intermediate calculations. Input your answer as a percent rounded to 2 decimal places.)

b. Compute Ke (required rate of return on common equity based on the dividend valuation model). (Do not round intermediate calculations. Input your answer as a percent rounded to 2 decimal places.)
Answer #1

a). Cost of common equity = Rf + [β * (Km - Rf)]

= 8% + [1.7 * (13% - 8%)] = 8% + [1.7 * 5%] = 8% + 8.5% = 16.5%

b). Cost of common equity = [D1 / P0] + g

= [$0.90 / $20] + 0.09 = 0.045 + 0.09 = 0.135, or 13.50%